OpenStudy (love_to_love_you):

What is the equation of the tangent line passing through the point (1, 3) of the graph of the function f(x) = x^2 + x + 1?

OpenStudy (freckles):

Are you using \[\lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h}=f'(x)\] or you using short cut methods ?

OpenStudy (love_to_love_you):

Both really

OpenStudy (freckles):

Use either the definition or the short cut methods to differentiate x^2+x+1

OpenStudy (freckles):

This will give you the slope formula... This means this will give you the slope of the tangent line at any x

OpenStudy (freckles):

a line can with point (a,b) and slope m can be written as y-b=m(x-a) so we already know (a,b) which is (1,3) we just need to find m which is f' evaluated at (1,3) or just x=1 since our function for f' will be in terms of x

OpenStudy (love_to_love_you):

m = (y-3)/(x-1)

OpenStudy (freckles):

and remember y is given as x^2+x+1 and you also forgetting to insert the limit as x approaches 1 there

OpenStudy (love_to_love_you):

Blah, my math is terrible and I'm not very good at limits, can we work through that?

OpenStudy (freckles):

do you know how to simplify x^2+x+1-3?

OpenStudy (love_to_love_you):

x^2+x-2 ?

OpenStudy (freckles):

yes do you know how to factor x^2+x-2?

OpenStudy (love_to_love_you):

(x+2)(x-1)

OpenStudy (freckles):

and what do you have on bottom there ?

OpenStudy (freckles):

you had (y-3)/(x-1) as x approaches 1

OpenStudy (freckles):

you factored y-3 as (x+2)(x-1)

OpenStudy (freckles):

so you have (x+2)(x-1)/(x-1) as x approaches 1

OpenStudy (freckles):

since you know x-1 isn't 0 you can say (x-1)/(x-1)=?

OpenStudy (freckles):

5/5=? 6/6=? -1/-1=? -8/-8=? ... (x-1)/(x-1)=? assuming x-1 isn't ever 0

OpenStudy (freckles):

something/(the same something)=1 when the something isn't zero

OpenStudy (love_to_love_you):

Sorry my tutor came over, it would be 1

OpenStudy (freckles):

yes so you are left with (x+2) as x approaches 1

OpenStudy (freckles):

\[y=x^2+x+1 \\ m=\lim_{x \rightarrow 1}\frac{y-3}{x-1}=\lim_{x \rightarrow 1} \frac{(x^2+x+1)-3}{x-1} \ \\ =\lim_{x \rightarrow 1}\frac{x^2+x-2}{x-1}=\lim_{x \rightarrow 1} \frac{(x+2)(x-1)}{x-1}=\lim_{x \rightarrow 1} (x+2)\]

OpenStudy (freckles):

do you know how to finish evaluating the slope there?